OpenStudy (anonymous):

Simiplify (25 x - power to the 5 ) power to 25

OpenStudy (campbell_st):

so its simplify \[(25x^5)^{25}\]

OpenStudy (anonymous):

yes

OpenStudy (campbell_st):

ok.... lets start by splitting things \[(25x^5)^{25} = (25)^{25}(x^5)^{25}\] the power of a power law says \[(m^a)^b = m^{a \times b}\] you can use the law here for x

OpenStudy (campbell_st):

I'm just unsure if \[25^{25} = (5^2)^{25} = 5^{50}\] of its left as it is

OpenStudy (anonymous):

Oh i understand how to do it now. thank you.

OpenStudy (campbell_st):

hope it helps